# Two-dimensional fluids via matrix hydrodynamics

### v3

@@ 1. Introduction
If you stir up an incompressible, low-viscosity fluid confined to a thin (almost 2-D) bounded domain then, under conditions on the initial configuration and the applied forcing, what you see is spectacular, completely different from the situation in full 3-D. The fluid self-organizes into large, coherently swirling vortex condensates. [^1] The effect can be experimentally reproduced by a thin soap film flowing rapidly through a fine comb [13], or by a conducting fluid confined to a thin layer and driven into turbulence by a temporally varying magnetic field [51]. Some regions swirl counterclockwise (positive vorticity) while others swirl clockwise (negative vorticity), with occasional merger between equal-signed regions, and repulsion between opposite-signed regions. What is the mechanism of vortex condensation and under which conditions does it occur? More mathematically formulated, what is the generic long-time behavior of the 2-D Euler equations?
@@
A milestone was reached in 1949, when Onsager [48] suggested to use statistical mechanics applied to a large but finite number $n$ of point vortices on a flat torus (doubly periodic square). Onsager ingeniously realized that if the energy of a configuration is large enough, the thermodynamical temperature is negative. Point vortices of equal sign will then tend to cluster “so as to use up excess energy at the least possible cost in terms of degrees of freedom” [48, p. 281]. Thus, statistical mechanics predicts vortex condensation. Under the ergodicity assumption, it also predicts the long-time states corresponding to “thermal equilibrium”. Onsager’s ideas have fostered several mathematically rigorous results, for example weak convergence, in both positive and negative temperature regimes, of the canonical Gibbs measure as $n\to\infty$ [7, 34].
@@
But there is a problem with the point vortex approach, as Onsager himself pointed out: “When we compare our idealized model with reality, we have to admit one profound difference: the distribution of vorticity which occur in the actual flow of normal liquids are continuous” [48, p. 281]. In particular, solutions to the 2-D Euler equations with continuous, bounded vorticity possess a rich geometric structure: they are described by an infinite-dimensional Lie–Poisson system, which gives rise to infinitely many conservation laws given by Casimir functions (cf. Sec. 2 below). Whereas point-vortices in principle adhere to this structure, with Casimirs corresponding to point vortex strengths, they do it in a weak sense which fails to capture the effects of the incompressibility for bounded vortices, for example. Furthermore, there is evidence that the continuous Casimirs significantly affect the long-time behavior (cf. Abramov and Majda [1]). A further criticism of Onsager’s statistical model is that the ergodicity assumption is not valid (which does not, however, wholly invalidate the statistical theory; see the discussion in Marchioro and Pulvirenti [42, Sec. 7.5]). From these perspectives, Onsager’s model is inadequate.
@@
Another established technique for 2-D hydrodynamics is non-linear PDE analysis. It has provided many rigorous and deep results, but at the cost of restricted settings, for example perturbations of steady solutions. We refer to the monograph of Marchioro and Pulvirenti [42] and the lecture notes of Šverák [53] for an overview. The grand vision, of course, is to bridge PDE analysis results with predictions from statistical hydrodynamics. It is a notoriously difficult problem. A flavor of the difficulty is captured in the following “incompatibility” between the two approaches: The statistically predicted vortex condensation implies merging of equally signed vorticity regions. Yet, in the 2-D Euler equations, vorticity is advected by the fluid velocity field, so merging can only occur by trapping increasingly narrow, but deep, vorticity variations in an overwhelmingly complicated stretch-and-fold process. These variations cannot “disappear” in the $C^{1}$-topology. Indeed, the $C^{1}$-norm along regular solutions can, and in numerical simulations typically does, grow extremely fast (e.g., double exponential [37]).
@@
A third tool for 2-D hydrodynamics is to carry out numerical experiments (e.g., Fig. 1, where the mixing phase (b) captures a typical stretch-and-fold process). Indeed, since the work of Lorentz [41] in the 1960s, computer simulations have guided many analytical results in fluid dynamics; recent examples are found in the references [14, 50, 37, 20, 9, 57]. But how can we know if the numerical results are consistent with the geometric structures principal to the long-time behavior? One approach is to ensure that the finite-dimensional dynamical model resulting from the discretization is itself a Lie–Poisson system, and that the finite-dimensional Lie–Poisson bracket approximates the infinite-dimensional one. Such a model was provided by Zeitlin [60, 61], based on quantization results of Hoppe [27, 28]. It is the only known Lie–Poisson spatial discretization of 2-D Euler, and numerical simulations indicate that it captures the qualitative behavior more accurately than traditional discretizations (e.g., the predicted spectral scaling law, which standard methods fail to catch even at significantly higher resolutions [10]). Yet, the model is surprisingly underexplored.
@@
Our paper draws from the following observation: in addition to numerical advantages, matrix hydrodynamics – the isospectral matrix flows enabled via Zeitlin’s model – brings theoretical insights to classical 2-D hydrodynamics, as it provides a link to matrix theory with all its findings. We therefore advocate matrix hydrodynamics as a tool for 2-D turbulence, similar in spirit to the point vortex model of Onsager, but with the benefit that it captures the Lie–Poisson geometry of continuous vorticity fields, including conservation of continuous Casimirs.
@@
In addition to the examples in this paper, a motivation and feasibility check for the approach comes from complex geometry. Indeed, in the late 1980s, quantization was promoted to address problems related to Kähler–Einstein metrics and the Calabi conjecture. Tian [54] then gave convergence results, since used regularly in theoretical contexts. Only much later did Donaldson [17] use quantization in a numerical study of Kähler–Einstein metrics. This is the antithesis to the developments in 2-D hydrodynamics, where Zeitlin’s model has been a numerical tool, until recently.
@@
The primary aim of the paper is to give basic convergence results for matrix hydrodynamics in the spherical setting (Theorem 4, Theorem 6, Corollary 17, and Theorem 19), and to showcase the notion in selected examples and relate it to existing theories (Sec. 4-6). The secondary aim, addressed in Sec. 3, is to give a transparent derivation of Zeitlin’s model on the sphere, based on unitary representation theory and focused on geometry rather than algebraic formulae. Before that, we review in Sec. 2 the geometric structure of the 2-D Euler equations.
@@
Acknowledgements This work was supported by the Swedish Research Council (grant number 2022-03453), the Knut and Alice Wallenberg Foundation (grant numbers WAF2019.0201 and KAW2020.0287), and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. The computations were enabled by resources provided by Chalmers e-Commons at Chalmers. The data handling was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. The authors would like to thank Michele Dolce and Theo Drivas for fruitful discussions related to this work, and Michael Roop for important comments related to the convergence proof in Appendix A.
@@ 2. Background: geometry of the 2-D Euler equations
For an incompressible, inviscid fluid, Euler [21] showed how Newton’s equations give rise to a set of partial differential equations (PDEs) for the velocity field $\boldsymbol{v}(x,t)$ of the motion,
@@
The sphere is a symplectic manifold for the symplectic form given by the spherical area form $\mu$ (in spherical coordinates $(\theta,\phi)\in[0,\pi]\times[0,2\pi)$ the area form is $\mu=\sin\theta\,\mathrm{d}\theta\wedge\mathrm{d}\phi$). Given a Hamiltonian function $f$ on $\mathbb{S}^{2}$, the corresponding symplectic vector field $X_{f}$ is defined by
@@
If we identify 1-forms with vector fields using the Riemannian structure of $\mathbb{S}^{2}$, equation (3) takes the form
@@
The space $\mathfrak{X}_{\mu}(\mathbb{S}^{2})$ of smooth, symplectic vector fields on $\mathbb{S}^{2}$ forms an infinite-dimensional Lie algebra (with the vector field bracket $\cdot,\cdot]_{\mathfrak{X}}$). This algebra is isomorphic to the Poisson algebra of smooth functions modulo constants $C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$ via the mapping $\psi\mapsto X_{\psi}$. Indeed, we have that
@@
Let us now describe the connection between this infinite-dimensional Lie algebra and the vorticity equation (2).
@@
In other words, the smooth dual of the Lie algebra of Hamiltonian functions (or equivalently symplectic vector fields) is given by smooth functions with vanishing mean.
@@
These were found by Helmholtz [25], and analyzed by Kirchhoff [35] and Kelvin [31]. Much later, Onsager [48] used them is his theory of statistical hydrodynamics. Let us see how they fit into the geometric framework of Lie-Poisson dynamics, following Marsden and Weinstein [43]. In general, if $G$ acts on a symplectic manifold $(M,\nu)$ by symplectomorphisms then there is a corresponding infinitesimal action vector field $\xi_{M}\in\mathfrak{X}_{\nu}(M)$ of $\xi\in\mathfrak{g}$. Suppose that $\xi_{M}$ is Hamiltonian, i.e., there is a function $H_{\xi}\in C_{0}^{\infty}(M)$ such that $\iota_{\xi_{M}}\nu=\mathrm{d}H_{\xi}$. Since $H_{\xi}$ is linear in $\xi$, we have that
$H_{\xi}(m)=\langle p(m),\xi\rangle$
for some function $p\colon M\to\mathfrak{g}^{*}$ called the momentum map. Usually (but not always) the momentum map is equivariant, i.e., $p(g\cdot m)=\mathrm{Ad}^{*}_{g}(p(m))$. If, in addition, the action of $G$ on $M$ is transitive then the image $p(M)\subset\mathfrak{g}^{*}$ is a (generally weak) coadjoint orbit.
@@
This is the transport map formulation of equation (6). It is the basis for Yudovich’s [59] global well-posedness results in $L^{\infty}$, which shows convergence of the fixed point iteration obtained by leap-frogging between the equations (8) for $\Phi$ (Picard iterations for ODE) and for $\psi$ (stationary PDE). Geometrically, the transport formulation captures that the symplectic leaves are generated by the action of the Lie group.
@@ 3.1. Quantization via representation
The result in Theorem 4 (and Theorem 19) is analogous to the $L^{2}$ convergence result of Gallagher [23] for Zeitlin’s original model on the flat torus $\mathbb{T}^{2}$. But there are technical differences in the proof: Gallagher’s result rely on the algebraic multiplication properties of the Fourier basis in $\mathbb{T}^{2}$, which are not available on $\mathbb{S}^{2}$.
In summary, we have a correspondence between functions and matrices that allows transcription of ideas and results from 2-D hydrodynamics to matrix theory. The dictionary in Table 1 contains some direct examples.
@@ 4. Coadjoint orbits and their closure
However, while solutions with smooth initial data remain smooth, the $C^{k}$-norms for $k\geq 1$ typically grow fast as the vorticity level-curves quickly become extremely entangled. This mechanism results in a Reeb graph that is well-defined but futile to track at moderate scales. It is therefore natural to study the closure of coadjoint orbits in a weaker topology, for instance $L^{\infty}$ as Yudovich’ existence theory suggests. In doing so, the Reeb graph is not stable in $L^{\infty}$ since it is based on critical values and Morse functions that require $C^{2}$ topology (see Fig. 3). Instead, what persists is the level-set measure $\lambda_{\omega}$, obtained by “flattening” the Reeb graph as follows. Recall the Casimir invariants $C_{f}(\omega)$. The mapping $C^{\infty}(\mathbb{R})\ni f\mapsto C_{f}(\omega)$ is a distribution with compact support, and $\lambda_{\omega}$ is the corresponding positive Radon measure on $\mathbb{R}$ (see Fig. 2b). The $L^{\infty}$ closure of a coadjoint orbit then fulfills $\overline{\mathcal{O}(\omega_{0})}\subseteq\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}$. (The latter set is characterized in Sec. 4.1 below.)
@@
The coadjoint action of $F\in\mathrm{SU}(N)$ on $W\in\mathfrak{su}(N)$ is $FWF^{\dagger}$. From the spectral theorem it then follows that the coadjoint orbit $\mathcal{O}^{N}(W)=\{FWF^{\dagger}\mid F\in\mathrm{SU}(N)\}$ consists of all elements in $\mathfrak{su}(N)$ with the same spectrum as $W$, or, equivalently, the same empirical measure $\lambda^{N}_{W}$. It is instructive to compare with the $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$ coadjoint orbits: in spirit, $\mathcal{O}^{N}(W)$ corresponds to $\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}$, situated between the smooth orbit $\mathcal{O}(\omega)$ (characterized by its measured Reeb graph) and Shnirelman’s weak closure $\overline{\mathcal{O}(\omega_{0})}^{*}$ (characterized by “convex” vortex mixing, cf. [16]).
For a fixed $N$, matrix hydrodynamics thereby behaves as the Euler equations in the strong $L^{\infty}$ topology: Casimirs are exactly conserved. But we can also regard all $N$ simultaneously, which gives a hierarchy of increasingly refined topologies.

### v4

@@ 1. Introduction
If you stir up an incompressible, low-viscosity fluid confined to a thin (almost 2-D) bounded domain, then, under conditions on the initial configuration and the applied forcing, what you see is spectacular, completely different from the situation in full 3-D. The fluid self-organizes into large, coherently swirling regions called vortex condensates. [^1] The effect can be experimentally reproduced by a thin soap film flowing rapidly through a fine comb [13], or by a conducting fluid confined to a thin layer and driven into turbulence by a temporally varying magnetic field [51]. Some regions swirl counterclockwise (positive vorticity), others clockwise (negative vorticity), with occasional merger between equal-signed regions, and repulsion between opposite-signed regions. What is the mechanism of vortex condensation and under which conditions does it occur? More mathematically formulated, what is the generic long-time behavior of the 2-D Euler equations?
@@
Towards an answer, a milestone was reached in 1949 when Onsager [48] applied statistical mechanics to a large but finite number of point vortices on a flat torus (doubly periodic square). Onsager understood that if the energy of a configuration is large, relative to the vortex strengths, then the thermodynamical temperature is negative. Point vortices of equal sign will then tend to cluster “so as to use up excess energy at the least possible cost in terms of degrees of freedom” [48, p. 281]. Statistical mechanics thus predicts vortex condensation. Under the ergodicity assumption, it also predicts long-time states corresponding to thermal equilibrium. Since then, Onsager’s ideas have fostered several rigorous results, for example, weak convergence of the Gibbs measure, in both positive and negative temperature regimes, as the number of point vortices increase [7, 34].
@@
But there is a problem with the point vortex approach, as Onsager himself pointed out: “When we compare our idealized model with reality, we have to admit one profound difference: the distribution of vorticity which occur in the actual flow of normal liquids are continuous” [48, p. 281]. In particular, solutions to the 2-D Euler equations with continuous, bounded vorticity possess a rich geometric structure: they make up an infinite-dimensional Lie–Poisson system, with infinitely many conservation laws given by Casimir functions (cf. Sec. 2 below). Whereas point-vortices formally fit into this structure, with Casimirs corresponding to the vortex strengths, they do it in a weak sense, which fails to capture the incompressible nature of the fluid, for example. In addition, point vortices cannot capture relations between the continuous Casimirs, and those relations affect the long-time behavior (cf. Abramov and Majda [1]). Moreover, the ergodicity assumption for Onsager’s statistical approach is invalid. [^2] This does not, however, wholly discredit the statistical theory; see the discussion in Marchioro and Pulvirenti [42, Sec. 7.5]. So, in summary, the point vortex model is insufficient to describe the long-time behavior of continuous solutions.
Another established technique for 2-D hydrodynamics is non-linear PDE analysis. It has led to many rigorous and deep results, but at the cost of restricted settings, for example perturbations of steady solutions. We refer to the monograph of Marchioro and Pulvirenti [42] and the lecture notes of Šverák [53] for an overview.
@@
The grand vision is to bridge PDE analysis results with predictions from statistical hydrodynamics. It is a notoriously difficult problem. As a flavor of the difficulty, the two approaches seem incompatible: The statistically predicted vortex condensation implies merging of equally signed vorticity regions. Yet, in the dynamics of the 2-D Euler equations, vorticity is advected by the fluid velocity field, so merging can only occur by trapping increasingly narrow, but deep, vorticity variations in an overwhelmingly complicated stretch-and-fold process. These variations cannot disappear in the $C^{1}$-topology. Indeed, the $C^{1}$-norm along regular solutions can grow extremely fast, up to double exponential [37], and in numerical simulations it typically does grow extremely fast.
@@
A third tool for 2-D hydrodynamics is to carry out numerical experiments (as illustrated in Fig. 1, where the mixing phase (b) captures a typical stretch-and-fold process). Indeed, since the work of Lorentz [41] in the 1960s, computer simulations have guided many analytical results in fluid dynamics; recent examples are found in the references [14, 50, 37, 20, 9, 57]. But are the numerical outcomes consistent with the geometry known to influence the long-time behavior?
To address this question, we can ask for the finite-dimensional dynamical model resulting from the discretization to be itself a Lie–Poisson system, whose finite-dimensional Lie–Poisson bracket approximates the infinite-dimensional one. Such a model was given by Zeitlin [60, 61], based on quantization results of Hoppe [27, 28]. It is the only known Lie–Poisson discretization of the 2-D Euler equations, and numerical simulations indicate that it captures the qualitative behavior better than traditional discretizations (e.g., the predicted spectral scaling law, which standard methods fail to catch even at significantly higher resolutions [10]). Yet, the model is vastly underexplored.
@@
This paper began with a simple idea: in addition to numerical advantages, matrix hydrodynamics – the matrix flows enabled via Zeitlin’s model – brings theoretical insights to classical 2-D hydrodynamics by providing a link to the rich and well-developed theory of matrices. The point is to put forward matrix hydrodynamics as a tool for 2-D turbulence, similar in spirit to the point vortex model of Onsager, but with the benefit of capturing the Lie–Poisson geometry of continuous vorticity fields, including conservation of continuous Casimirs.
@@
Besides the examples we give throughout the paper, a motivation and feasibility check for our idea comes from complex geometry. Indeed, in the late 1980s, quantization was promoted to address questions about Kähler–Einstein metrics and the Calabi conjecture. Tian [54] then gave convergence results, since used regularly in theoretical contexts. Only much later did Donaldson [17] use quantization in a numerical study of Kähler–Einstein metrics; the antithesis to the developments in 2-D hydrodynamics, where Zeitlin’s model has been a numerical tool, until recently.
@@
Our primary aim with the paper is to give basic convergence results for matrix hydrodynamics in the spherical setting (Theorem 4, Theorem 6, Corollary 17, and Theorem 19), and to showcase the fluid-to-matrix link in selected examples and relate it to existing theories (Sec. 4-6). Our secondary aim, addressed in Sec. 3, is to give a transparent derivation of Zeitlin’s model on the sphere, based on unitary representation theory and focused on geometry rather than algebraic formulae. But first, we review in Sec. 2 the geometry of the 2-D Euler equations.
@@
Acknowledgements This work was supported by the Swedish Research Council (grant number 2022-03453), the Knut and Alice Wallenberg Foundation (grant numbers WAF2019.0201 and KAW2020.0287), and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. The computations were enabled by resources provided by Chalmers e-Commons at Chalmers. The data handling was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We would like to thank Michele Dolce and Theo Drivas for fruitful discussions related to this work, and Michael Roop and Ali Suri for important comments related to the convergence proof in Appendix A.
@@ 2. Background: geometry of the 2-D Euler equations
For an incompressible, inviscid fluid, Euler [21] showed how Newton’s second law leads to partial differential equations (PDEs) for the velocity field $\boldsymbol{v}(x,t)$ of the fluid’s motion,
@@
The sphere is a symplectic manifold with symplectic form given by the spherical area form $\mu$ (in spherical coordinates $(\theta,\phi)\in[0,\pi]\times[0,2\pi)$ the area form is $\mu=\sin\theta\,\mathrm{d}\theta\wedge\mathrm{d}\phi$). Given a Hamiltonian function $f$ on $\mathbb{S}^{2}$, the corresponding symplectic vector field $X_{f}$ is defined by
@@
If we identify 1-forms with vector fields using the Riemannian structure of $\mathbb{S}^{2}$, equation (3) becomes
@@
The space $\mathfrak{X}_{\mu}(\mathbb{S}^{2})$ of smooth, symplectic vector fields on $\mathbb{S}^{2}$ forms an infinite-dimensional Lie algebra (with the vector field bracket $\cdot,\cdot]_{\mathfrak{X}}$). This algebra is isomorphic to the Poisson algebra of smooth functions modulo constants $C^{\infty}(\mathbb{S}^{2})/\mathbb{R}$ via the mapping $\psi\mapsto X_{\psi}$. Indeed,
@@
Let us now describe the connection between the infinite-dimensional Lie algebra $\mathfrak{X}_{\mu}(\mathbb{S}^{2})$ and the vorticity equation (2).
@@
So the smooth dual of the Lie algebra of Hamiltonian functions (or equivalently symplectic vector fields) is given by smooth functions with vanishing mean.
@@
These were found by Helmholtz [25], and analyzed by Kirchhoff [35] and Kelvin [31]. Much later, Onsager [48] used them is his theory of statistical hydrodynamics. Let us see how they fit into the geometric framework of Lie-Poisson dynamics, following Marsden and Weinstein [43]. In general, if $G$ acts on a symplectic manifold $(M,\nu)$ by symplectomorphisms then there is a corresponding infinitesimal action vector field $\xi_{M}\in\mathfrak{X}_{\nu}(M)$ of $\xi\in\mathfrak{g}$. Suppose that $\xi_{M}$ is Hamiltonian, i.e., there is a function $H_{\xi}\in C_{0}^{\infty}(M)$ such that $\iota_{\xi_{M}}\nu=\mathrm{d}H_{\xi}$. Notice that $H_{\xi}$ must be linear in $\xi$. Suppose moreover there is a smooth function $p\colon M\to\mathfrak{g}^{*}$ such that
$H_{\xi}(m)=\langle p(m),\xi\rangle.$
Then $p$ is a momentum map. Often (but not always) we can take the momentum map to be equivariant, i.e., $p(g\cdot m)=\mathrm{Ad}^{*}_{g}(p(m))$. If, in addition, the action of $G$ on $M$ is transitive, then the image $p(M)\subset\mathfrak{g}^{*}$ is a coadjoint orbit (generally weak).
@@
This is the transport map formulation of equation (6). It is the basis for global well-posedness results (cf. [59, 42]), which shows convergence of the fixed point iteration obtained by leap-frogging between the equations (8) for $\Phi$ (Picard iterations for ODE) and for $\psi$ (stationary PDE). Geometrically, the transport formulation captures that the symplectic leaves are generated by the action of the Lie group.
@@ 3.1. Quantization via representation
The result in Theorem 4 (and Theorem 19) is analogous to the $L^{2}$ convergence result of Gallagher [23] for Zeitlin’s original model on the flat torus $\mathbb{T}^{2}$. But there are technical differences in the proof: Gallagher’s result rely on algebraic multiplication properties of the Fourier basis in $\mathbb{T}^{2}$, not available on $\mathbb{S}^{2}$.
In summary, we have a correspondence between functions and matrices that allows transcription of ideas and results from 2-D hydrodynamics to matrix theory. The dictionary in Table 1 contains the most direct examples.
@@ 4. Coadjoint orbits and their closure
However, while solutions with smooth initial data remain smooth, the $C^{k}$-norms for $k\geq 1$ grow fast as the vorticity level-curves quickly become extremely entangled. This mechanism results in a Reeb graph that is well-defined but futile to track at moderate scales. It is therefore natural to study the closure of coadjoint orbits in a weaker topology, for instance $L^{\infty}$ as Yudovich’ existence theory suggests. But the Reeb graph is not stable in $L^{\infty}$ since it is based on critical values and Morse functions that require $C^{2}$ topology (see Fig. 3). Instead, what persists is the level-set measure $\lambda_{\omega}$, obtained by “flattening” the Reeb graph as follows. Recall the Casimir invariants $C_{f}(\omega)$. The mapping $C^{\infty}(\mathbb{R})\ni f\mapsto C_{f}(\omega)$ is a distribution with compact support, and $\lambda_{\omega}$ is the corresponding positive Radon measure on $\mathbb{R}$ (see Fig. 2b). The $L^{\infty}$ closure of a coadjoint orbit then fulfills $\overline{\mathcal{O}(\omega_{0})}\subseteq\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}$. (The latter set is characterized in Sec. 4.1 below.)
@@
Since the coadjoint action of $F\in\mathrm{SU}(N)$ on $W\in\mathfrak{su}(N)$ is $FWF^{\dagger}$, it follows from the spectral theorem that the coadjoint orbit $\mathcal{O}^{N}(W)=\{FWF^{\dagger}\mid F\in\mathrm{SU}(N)\}$ is the set of skew-Hermitian matrices with the same spectrum as $W$, or, equivalently, the same empirical measure $\lambda^{N}_{W}$.
It is instructive to compare with coadjoint orbits for $\operatorname{Diff}_{\mu}(\mathbb{S}^{2})$: the finite-dimensional coadjoint orbit $\mathcal{O}^{N}(W)$ corresponds (see Proposition 11 below) to the set
$\overline{\mathcal{O}}(\omega_{0})\coloneqq\{\omega\in L^{\infty}(\mathbb{S}^{2})\mid\lambda_{\omega}=\lambda_{\omega_{0}}\}.$
This set is strictly larger than the smooth orbit $\mathcal{O}(\omega_{0})$ (characterized by its measured Reeb graph), but strictly smaller than Shnirelman’s weak closure $\overline{\mathcal{O}(\omega_{0})}^{*}$ (characterized by convex vortex mixing, cf. [16]). For a fixed $N$, matrix hydrodynamics thereby behaves as the Euler equations in the strong $L^{\infty}$ topology: Casimirs are exactly conserved. But we can also regard all $N$ simultaneously, which gives a hierarchy of increasingly refined topologies.